

********************************************************************
UARS WEEKLY STATUS REPORT - 04 AUG 2000
********************************************************************
***********
SOLSTICE
***********
SOLSTICE is in good health, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. The solar data is not affected by the larger than normal
pointing errors since they are within our calibrations. Additionally, we
continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
Level2 processing is halted due to bugs in the OA_SSPP_ATT service. CDHF
and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM continues to gather a normal set of daily solar scans scheduled
for periods of predicted telemetry. Solar pointing quality has become
a concern due to the loss of the star tracker. We are evaluating the
impact in more detail.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM/AXIS-1 was activated on July 27 and deactivated on July 22. The
full PEM/AXIS was activated July 24 and deactivated July 26. The PEM
zenith spectrometers were reactivated in daylight mode on July 21. All of
PEM was off for the July 12 particle event.
At the present time, PEM energy deposition is not being generated.
Recent data configuration changes have affected PEM software causing multiple
failures. These failures are being reviewed by the PEM team. In order to
correct for new telemetry requirements, PEM will need to update processing
code with a patch delivery. The delivery patch will not include changes which
would allow the PEM software to execute on a DEC Alpha machine.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE was powered on end of day August 1, 2000. All engineering parameters
are within nominal operating region. Science data channel exoatmospheric
values are nominal.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight Operations
Team and to assist the Project Scientist in scheduling of instrument
operations.
The MPG attended a meeting to discuss Fixed Head Star Tracker 1 on
Wednesday, August 2, 2000. The MPG was asked to query the instrument
teams about their instrument ability or resources available to be used
for attitude determination in case FHST1 becomes unusable.
We are waiting for the answers.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
3 days SEN (5.0) PROD (1992) (3R-3C=0L)
UARS CDHF REHOST:
- Production Control has started parallel testing Monday, 31-Jul-2000.
This testing on the Rehost System will be against all active instruments
in the Test Mode. Same UARS data will be processed on the current system
in Production Mode. PI Representatives will be notified regarding this
plan.
- UARS CDHF Rehost Operational Readiness Review was held August 3,2000.
A tour of the facility was given after the presentation.
***********
FOT
***********
REPORTING PERIOD
This report covers 22 July 2000 (Orbit 48437, GMT Day 204) through 28 July 2000
(Orbit 48542 , GMT Day 210).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
AXIS-1 204/2353 48451 Off
PEM 206/1635 48477 AXIS On
AXIS-2 208/0520 48500 Off
AXIS-1 208/0757 48502 Off
The beta angle ranged between the angles of 67.2 degrees to 71.3 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+2 degrees on 14 August 2000 (DOY 227).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.5 to 0.8 degrees C. Additional battery performance data for each
battery on 21 July 2000 and 28 June 2000 are provided in Appendix D.
The clock error ranged from -10.6 to 10.6 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 00
% ALERTS = 02
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 07
% PEM = 04
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 02
Two (2) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/145 - 2287.497008
- Transponder B: 00/152 - 2287.498920
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/202 show the observatory data loss to be 296 hours, 44
minutes, 04.705 seconds (an increase of 5 minute 40.968 seconds since the last
report period). This is a 0.40 percent data loss which equals a 99.60 percent
data capture for the mission.
The increase of 5 minute 40.968 seconds occurred over four (4) days:
- 0 minutes 07.368 seconds on Day 204 due to a data hit,
- 2 minutes 26.880 seconds on Day 206 due to late acquisition,
- 0 minutes 10.272 seconds on Day 207 due to a data hit,
- 2 minutes 56.448 seconds on Day 208 due to a late acquisition.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS continues to operate FHST-1 this week. The alignment matrix loaded on 21
July 2000 was not accurate so the existing alignment matrix was put back in.
Some better guesses to the alignment were loaded on 26 and 28 July 2000,
according to calculations by Dave Lorenz, until a more accurate alignment can
be given by FDF.
The FOT has delayed the testing of the new TMONs until the FHST-2 anomaly is
resolved. No time frame has been set as to when testing will resume.
The FOT has begun writing the online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure will
be tested on the simulator and put online for immediate execution should an
anomalous condition take place.
MEETINGS
The FOT attended a meeting to discuss FHST-1 on 26 July 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for August 14, 2000 (DOY 227). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 28 July 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
48439 204/0430 00-361 NBTR PLAYBACK
48451 204/2353 00-363 AXIS-1 Power OFF
48463 205/1812 00-364 OBC Clk Adj fr 41.2 to 20.6 msec/day
48477 206/1635 00-365 PEM Power ON PEM AXIS 1 & 2
48479 206/1954 00-366 ACS FHST1 RECAL
48487 207/0938 00-367 OBC Clk Adj 20.6 to 41.2 msec/day
48500 208/0520 00-368 AXIS-2 Power OFF
48502 208/0757 00-369 AXIS-1 Power OFF
48509 208/2040 00-370 FHST-1 CALIBRATION
48522 209/1723 00-371 OBC Clk Adj fr 41.6 to 20.6 msec/day
48536 210/1502 00-372 FHST1 ROTATIONAL ALIGNMENT
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 28 July 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
OBSZ1OOL UFLTR Z1 Residual Exceeded 100 Arcsec 48470, 48471,7,8,9
OBSZ2OOL UFLTR Z2 Residual Exceeded 100 Arcsec 48536-8
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 28 July 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
48463 205/18:12 (23 July 00) from 41.2 to 20.6 msec/day
48487 207/09:51 (25 July 00) from 20.6 to 41.2 msec/day
48522 209/17:28 (27 July 00) from 41.6 to 20.6 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 28 July 2000
21 July 2000 - GMT Day 203 Beta = 67.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------ ------- ------
1 +56.0/+39.2 -3.05/-4.62 +0.0/+0.0 15.5 0.000 0.0
2 +16.8/-11.2 +3.55/+1.97 +19.2/-15.2 26.1 1.040 17.4
3 +22.4/-11.2 +2.60/+1.34 +19.2/-15.2 25.9 1.030 17.1
28 July 2000 - GMT Day 210 Beta = 71.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------ ------- ------
1 +67.2/+44.8 -2.42/-3.68 +0.0/+0.0 15.5 0.000 0.0
2 +16.8/-11.2 +3.23/+1.66 +17.6/-13.6 27.0 1.028 14.8
3 +22.4/-22.4 +2.60/+1.34 +16.8/-13.6 27.0 1.030 14.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 28 July 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-031 FOT 48469 206/0434 14170 Generic Late Acq.
00-033 FOT 48509 208/1958 2723 Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-026 FOT see list June 00 14170 Generic Late Acq.
00-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
SEVNTFW00/30